how do I go about solving these linear equations using substitution
4x-y=5
4x+4y=-4
I know I have to add to both sides of the first one but I do not remember which sides or where to go from there.

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

how do I go about solving these linear equations using substitution
4x-y=5 --> y = 4x-5
4x+4y=-4
-----
Sub for y in the 2nd eqn
4x + 4(4x-5) = -4
x + 4x-5 = -1
5x = 4
x = 4/5
------
y = -9/5

Answer by stanbon(75887)

go about solving these linear equations using substitution
4x-y=5
4x+4y=-4
----
Modify::
y = 4x-5
x + y = -1
-------------------
Substitute for "y" and solve for "x"::
x + 4x-5 = -1
----
5x = 4
x = 4/5
----
Solve for "y"
x + y = -1
4/5 + y = -1
y = -9/5
--------------
Cheers,
Stan H.
--------------